OpenStudy (driftracer305):

In flipping a coin, heads and tails are mutually exclusive events (a coin can't show both heads and tails at the same time). If two coins are flipped, the outcomes may be heads-heads, heads-tails, tails-heads, or tails-tails. However, tails plus heads can occur in two ways. Therefore the probability of getting a tails-plus-heads combination is.....?

jimthompson5910 (jim_thompson5910):

Hint: The sample space looks like this HH HT TH TT where H stands for heads and T stands for tails

OpenStudy (driftracer305):

right

jimthompson5910 (jim_thompson5910):

So count the number of ways to get tails and heads together. Then divide that by 4 (since there are 4 total possible combos)

OpenStudy (driftracer305):

im just confused.....does this take to the product or sum rule or something totally different?

OpenStudy (driftracer305):

ok so that is 1/2...

jimthompson5910 (jim_thompson5910):

yes 2/4 = 1/2

jimthompson5910 (jim_thompson5910):

I'm not sure about the product or sum rule part you asked about. It doesn't seem like you need either rule.

OpenStudy (driftracer305):

gotcha

OpenStudy (driftracer305):

so would that mean 1/4 + 1/4 = 1/2 ?

jimthompson5910 (jim_thompson5910):

yeah I guess you can think of it like that 1/4 + 1/4 = (1+1)/4 = 2/4 = 1/2

OpenStudy (driftracer305):

right becoz one event HT occurs 1/4 and the other... TH also 1/4.....but i dont completely understand why it is a sum rule here then....

OpenStudy (driftracer305):

Thanks @jim_thompson5910 !

OpenStudy (shkh.zain):

Addition rule is applied when there is a word "or" used. E.g. Getting a head or tail